Which of the following calculations should be performed to obtain a “Total Variance” proof total?

Budgeted cost less applied cost

Actual cost less applied cost

Actual cost less budgeted cost

None of these is correct.
Answer #1

Answer: 3rd option

This is the difference of actual cost and budgeted cost.

Suppose in case of material variance, total material variance is proved as below:

[SP = standard or budgeted price; SQ = standard of budgeted quantity; AP = actual price; AQ = actual quantity]

Material price variance = SP × AQ – AP × AQ

Material quantity variance = SQ × SP – SP × AQ

Total material variance = Material price variance + Material quantity variance

                                    = [SP × AQ – AP × AQ] + [SQ × SP – SP × AQ]

                                    = SP × AQ – AP × AQ + SQ × SP – SP × AQ

                                    = SQ × SP - AP × AQ

                                    = Budgeted cost – Actual cost